OpenStudy (anonymous):

it was once recorded that a jaguar left skid marks with were 290 m in lenght. assuming that the jaguar skidded to a stop with a constant acceleration of -3.90 m/s^2, determine the speed of the jaguar before it began to skid.

OpenStudy (anonymous):

Here's a graphical solution. If you plot speed vs. time, and you have constant acceleration, you get a graph similar to [skid.png]. The distance traveled during the deceleration is the area beneath the line. In this case, it's a triangle. Therefore: \[d=\frac{1}{2}st\] And you know the acceleration and distance traveled: \[a=\frac{s}{t}=-3.90\] \[d=290\] We solve for \(t\): \[t=\frac{s}{-3.90}\] Substitute \(t\) to the first equation: \[\begin{align*} d&=\frac{1}{2}st\\ 290&=\frac{1}{2}\left( \frac{s^2}{-3.90} \right) \end{align*}\] We can now solve for initial speed \(s\), disregard the negative sign of the acceleration: \[s=\sqrt{(2)(290)(3.90)}\]